Find the value of "a" such that is a perfect square
y^2+2y+a

Answer by venugopalramana(3286)

Find the value of "a" such that is a perfect square
y^2+2y+a
USE (A+B)^2=A^2+2AB+B^2
(Y)^2+2(Y)(1)+(1)^2=(Y+1)^2...HENCE A=1